# AttentionPredictor: Temporal Pattern Matters for Efficient LLM Inference

### v2

@@ 6 Conclusion
We present AttentionPredictor, the first learning-based critical token identification approach for KV cache compression. With a lightweight convolution model, AttentionPredictor captures spatiotemporal patterns of attention score and predicts the next-token attention score accurately. Additionally, we propose the first cross-token prefetching framework, which effectively mitigates prediction and transfer delays in the LLM decoding stage. Experiments on the long-context datasets demonstrate that AttentionPredictor achieves comparable accuracy of LLM with 16× KV cache compression.
Appendix A Query Similarity
We evaluate the similarity of consecutive queries. Specifically, we extract query data from LLaMA-3.1-8B using LongBench and compute the cosine autocorrelation of adjacent queries with a lag of 1. As shown in Figure 6, the heatmap illustrates the query similarity across each layer and head of the LLM. The average similarity is 86% for Longchat and 87% for LLaMA-3.1, indicating strong continuity. Furthermore, the similarity in LLaMA is more consistent across layers, while Longchat shows higher similarity in the shallower layers.
@@ Appendix A Query Similarity
Appendix B Implementaion of AttentionPredictor
@@
Training Details. We capture the attention scores during full-cache inference as raw data. We package the attention scores of $H$ steps, together with the $(H+1)$-th step’s score, as input-output pairs. The training data is block-compressed to align with the usage during LLM model inference. Specifically, the input history attention is $A_{H}\in\mathbb{R}^{H\times\frac{t}{b}}$ and the output attention score is $S\in\mathbb{R}^{1\times\frac{t}{b}}$. Attention scores with insufficient lengths are padded with zeros at the end to ensure that all data within a package have the same length. For the attention at step $t+1$, we only predict the first $t$ positions because the last position is derived from the newly generated key, which does not need to be compressed or transferred. Notably, only the attention scores from the decoding phase are used as output data. As a result, the training dataset includes the final $H$ steps of the pre-filling attention and all attention scores from the decoding stage. To train the model, we calculate the discrepancy between the predicted and real attention scores using mean squared error (MSE) loss. The training epoch is set to 30, and we select the model with the best attention recovery rate as shown in Equation 4.
@@
Cross-token Prefetching. We leverage GPU parallel streams and CPU multi-threading to parallelize the token retrieval and cache transfer across different layers. Consistent with the main experiment, we maintain uncompressed caches for the first two layers to ensure higher inference accuracy.
Appendix C Experiments
C.1 Implementation of Baselines
@@
Our experiments are based on LLMs using 16-bit floating points, and we utilize FlashAttention2 (Dao, 2024) during the prefilling stage to reduce GPU memory consumption. Since FlashAttention2 does not output intermediate attention scores, for methods that rely on historical attention, we additionally compute partial prefill attention to obtain the scores. Our batch size is fixed at 1. For all baselines, we use the official code implementations to ensure the effectiveness of the methods. Additionally, for all methods, we skip compress the first two layers of the LLM.
@@
  • StreamingLLM (Xiao et al., 2024) finds that the sink token is very important, so it retains the initial and the most recent few tokens. For StreamingLLM, we evenly distribute the budget between sink tokens and local tokens.
@@
  • H2O (Zhang et al., 2023) accumulates all historical attention scores as an estimation. We evenly distribute the budget between the heavy budget and the recent budget. Since computing the full historical data can cause OOM with longer inputs, we use the last 64 steps of historical information to calculate the heavy hitters, and denote this as H2O+. Note that accumulating all prefill attention scores causes H2O to focus on initial tokens, so this is an improvement to the method’s effectiveness.
@@
  • SnapKV (Li et al., 2024b) accumulates attention scores within a recent window as an estimate and performs one-time filtering during the prefill stage. We set the time window to a size of 64 to align with our method. Since SnapKV only performs cache compression once during the prefill stage, the number of KV tokens it uses continues to grow during decoding, resulting in a relatively lower compression ratio compared to other methods.
@@
  • Quest (Tang et al., 2024) uses the current query and paged key to approximate the attention score. We use a chunk size of 16, corresponding to the block size in our method, which is also the parameter value used in the original paper.
C.2 Details of LongBench Dataset
@@
LongBench (Bai et al., 2024) is a carefully crafted benchmark suite designed to evaluate the capabilities of language models in processing extended documents and complex information sequences. It was developed for multi-task assessment of long-context inputs. The details of the metrics, number of words, language, and data used in LongBench are presented in Table 6.
@@
C.3 Construction of Long COT Dataset
To simulate Chain-of-Thought (CoT) tasks within long-context, we increased the number of few-shot examples. Specifically, we randomly selected a fixed number of questions and standard CoT answer pairs as prompts, along with the questions to be tested. We chose 25, 47, and 97 few-shot examples, resulting in input lengths of approximately 4K, 8K, and 16K tokens respectively. The few-shot data were sourced from the GSM8K training set. Since the test set does not overlap with the training set, the answers remain undisclosed to the LLM.
C.4 Longbench Results of All Tasks
Table 7: The detailed evaluation results from LongBench of all tasks.
@@ C.4 Longbench Results of All Tasks
4096 AttentionPredictor 28.13 45.48 55.81 56.48 49.41 30.46 33.57 24.77 27.10 72.50 90.26 41.45 9.32 99.50 55.88 46.53 47.92
C.5 Ablation Study Results
@@
Hyperparameter results. We evaluate the impact of hyperparameters on our method. Specifically, we train the predictor using the default parameters of the main experiment, i.e., $b=16$, $H=64$. We employ full attention without sparse computation during training, which is equivalent to performing distribution error calibration at each step, i.e., $M=1$. Subsequently, we test the performance under different hyperparameters using the trained model. Other parameters during testing remain consistent with the main experiment, with a KV budget of 1K. We evaluate six representative tasks in Longbench. The results are shown in Table 8.
When the block size is 8, we achieve comparable performance to the non-compressed model, demonstrating the effectiveness of our method. Additionally, we significantly mitigate the accuracy drop caused by the increase in retrieval granularity. When $b=64$, AttentionPredictor outperforms Quest by 10%. For the summary long output task, increasing the frequency of error calibration effectively improves model performance, underscoring the effectiveness of our error calibration method.

### v3

@@ 5 Conclusion
We present AttentionPredictor, the first learning-based method to directly predict attention patterns for KV compression and attention sparsity method. With a lightweight convolution model, AttentionPredictor captures spatiotemporal patterns of attention score and predicts the next-token attention score accurately. Additionally, we propose the first cross-token prefetching framework, which effectively mitigates prediction and transfer delays in the LLM decoding stage. Experiments on the long-context datasets demonstrate that AttentionPredictor achieves 13× KV cache compression and 5.6× speedup with comparable accuracy of LLM.
6 Acknowledgments
This work was supported in part by National Key R&D Program of China under contract 2022ZD0119801, National Nature Science Foundations of China grants U23A20388, 62021001, U19B2026, and U19B2044. This work was supported in part by Huawei as well. We would like to thank all the anonymous reviewers for their insightful comments.
Appendix A High Similarity of Queries and Keys
A.1 Qeury Similarity
Our model predicts the attention at step $t+1$ from a recent history. Since the temporal dimension corresponds to the query sequence (one query per step), the similarity between adjacent queries ($q_{t}$ and $q_{t+1}$) is the most critical factor for demonstrating step-by-step temporal continuity. Therefore, we evaluate the similarity of consecutive queries. Specifically, we extract query data from LLM using LongBench and compute the cosine autocorrelation of adjacent queries with a lag of 1:
$\rho(1)=\frac{1}{t-1}\sum_{i=1}^{t-1}\frac{\langle q_{i},q_{i+1}\rangle}{\|q_{i}\|\cdot\|q_{i+1}\|},$
As shown in Figure 9, the heatmap illustrates the query similarity across each layer and head of the LLM. The average similarity is 86% for Longchat and 87% for LLaMA-3.1, indicating strong continuity. Furthermore, the similarity in LLaMA is more consistent across layers, while Longchat shows higher similarity in the shallower layers.
@@ A.1 Qeury Similarity
Table 4: Self similarity of query and key over different distances.
Distance 1 2 5 10 20 50
Query Similarity 0.87 0.83 0.80 0.78 0.77 0.76
Key Similarity 0.82 0.77 0.74 0.72 0.71 0.67
To provide a more detailed analysis of the similarity of queries at a far distance, we measure the average query cosine similarity across various distances (lags). The results are presented in Table 4. The query similarity gradually decreases as the distance increases, which is expected. However, the similarity remains relatively high even at longer distances. This finding suggests that while the influence of the most recent tokens is strongest, the attention patterns exhibit a strong local stability that decays slowly over time. This supports our model’s effectiveness in capturing these evolving yet persistent patterns using a recent history window.
A.2 Key Similarity
In addition to queries, we also investigate the similarity of keys over time. When analyzing the attention scores between pairs such as $A_{t,i}$ and $A_{t+1,i}$, both attend to the same position $i$. Consequently, the keys involved in the two computations are exactly the same ($k_{i}$), and therefore no issue of key self-similarity arises in this case.
In contrast, when analyzing patterns like the "Periodic Slashes" pattern, the comparison involves pairs such as $A_{t,i}$ and $A_{t+1,i+1}$, where the attention scores are influenced by different keys ($k_{i}$ and $k_{i+1}$). In this case, the self-similarity of keys becomes relevant. To quantify this, we compute the cosine similarity between keys at different positions with various lags:
$\rho(\tau)=\frac{1}{t-\tau}\sum_{i=1}^{t-\tau}\frac{\langle k_{i},k_{i+\tau}\rangle}{\|k_{i}\|\cdot\|k_{i+\tau}\|},\quad\tau\geq 1,$
where $\tau$ denotes the distance between keys. The results are also presented in Table 4, and we can find that they indeed exhibit strong self-similarity. These consistently high similarity values indicate that keys vary slowly over time. Our findings are also consistent with prior works such as KIVI [34] and KVQuant [35], which reported that keys often contain fixed, massive channels, suggesting limited variation across positions.
Appendix B Implement Details
B.1 Implement Details of AttentionPredictor
Model Architecture Details. The architecture of the CNN-based AttentionPredictor model is summarized in Table 5. This design ensures efficiency and adaptability to varying sequence lengths, while maintaining sufficient capacity to capture temporal patterns.
Table 5: CNN-based AttentionPredictor Architecture.
Layer No. Layer Type Details
1 2D Convolution Kernel: (3,3), Padding: 1, Channels: 1 $\rightarrow$ 16
2 ReLU Activation -
3 2D Convolution Kernel: (3,3), Padding: 1, Channels: 16 $\rightarrow$ 32
4 ReLU Activation -
5 Adaptive Avg. Pooling Collapses temporal history dimension to size 1
6 1D Convolution Kernel: 1, Channels: 32 $\rightarrow$ 1
@@ 4 ReLU Activation -
Training Details. We capture the attention scores during full-cache inference as raw data. We package the attention scores of $H$ steps, together with the $(H+1)$-th step’s score, as input-output pairs. The training data is block-compressed to align with the usage during LLM model inference. Specifically, the input history attention is $\mathcal{A}_{H}\in\mathbb{R}^{H\times\frac{t}{b}}$ padding with zero for short attentions and the output attention score is $C\in\mathbb{R}^{1\times\frac{t}{b}}$. Attention scores with insufficient lengths are padded with zeros at the end to ensure that all data within a package has the same length. For the attention at step $t+1$, we only predict the first $t$ positions because the last position is derived from the newly generated key, which does not need to be compressed or transferred. Notably, only the attention scores from the decoding phase are used as output data. As a result, the training dataset includes the final $H$ steps of the pre-filling attention and all attention scores from the decoding stage. To train the model, we calculate the discrepancy between the predicted and real attention scores using mean squared error (MSE) loss. The training epoch is set to 30, and we select the model with the best attention prediction accuracy as shown in Equation 1.
@@
Cross-token Prefetching. We leverage GPU parallel streams and CPU multi-threading to parallelize the token retrieval and cache transfer across different layers. Consistent with the main experiment, we maintain uncompressed caches for the first two layers to ensure higher inference accuracy. The timeline of prefetching framework is in Figure 10.
B.2 Implement Details of Baselines
@@
Our experiments are based on LLMs using 16-bit floating points, and we utilize FlashAttention2 [56] during the prefilling stage to reduce GPU memory consumption. Since FlashAttention2 does not output intermediate attention scores, for methods that rely on historical attention, we additionally compute partial prefill attention to obtain the scores. Our batch size is fixed at 1. For all baselines, we use the official code implementations to ensure the effectiveness of the methods. Additionally, for all methods, we skip compress the first two layers of the LLM.
@@
  • StreamingLLM [12] finds that the sink token is very important, so it retains the initial and the most recent few tokens. For StreamingLLM, we evenly distribute the budget between sink tokens and local tokens.
@@
  • H2O [11] accumulates all historical attention scores as an estimation. We evenly distribute the budget between the heavy budget and the recent budget. Since computing the full historical data can cause OOM with longer inputs, we use the last 64 steps of historical information to calculate the heavy hitters, and denote this as H2O+. Note that accumulating all prefill attention scores causes H2O to focus on initial tokens, so this is an improvement to the method’s effectiveness.
@@
  • SnapKV [13] accumulates attention scores within a recent window as an estimate and performs one-time filtering during the prefill stage. We set the time window to a size of 64 to align with our method. Since SnapKV only performs cache compression once during the prefill stage, the number of KV tokens it uses continues to grow during decoding, resulting in a relatively lower compression ratio compared to other methods.
@@
  • Quest [31] uses the current query and paged key to approximate the attention score. We use a chunk size of 16, corresponding to the block size in our method, which is also the parameter value used in the original paper.
B.3 Details of LongBench Dataset
@@
LongBench [45] is a carefully crafted benchmark suite designed to evaluate the capabilities of language models in processing extended documents and complex information sequences. It was developed for multi-task assessment of long-context inputs. The details of the metrics, number of words, language, and data used in LongBench are presented in Table 6.
@@
B.4 Details of Efficiency Experiment Setup
Our efficiency evaluation is conducted under the cache offloading scenario, which is essential for long-context generation when the KV cache size exceeds the available GPU memory. Unlike GPU-based eviction/compression, offloading keeps the full cache on CPU, preserving model accuracy. This setting is also adopted in recent works [32, 19, 33].
We conduct experiments under this offloading setup on FlashAttention2, H2O, and our proposed AttentionPredictor:
  • FlashAttention2 [56]: We utilize the standard KV cache offloading implementation provided by the Hugging Face library. In this setup, at each decoding layer, the entire KV cache for the upcoming layers is prefetched from CPU to GPU to compute the full attention.
  • H2O [11]: Also uses cross-layer prefetching but transfers sparse KV cache, making the comparison a more direct evaluation of the prefetching framework’s efficiency.
  • AttentionPredictor: Runs on an offloading setup but prefetches across tokens, improving efficiency.
Appendix C Additional Experimental Results on Accuracy
C.1 Results on Long N-shot CoT GSM8K.
We evaluate the reasoning task under a long-input scenario, which is different with long output scenario in Section 4.2. As shown in Table 7, we evaluate our method on the mathematical reasoning dataset GSM8K with LLaMA-3.1. To simulate Chain-of-Thought (CoT) tasks within long-context, we increased the number of few-shot examples. Specifically, we randomly selected a fixed number of questions and standard CoT answer pairs as prompts, along with the questions to be tested. We chose 25, 47, and 97 few-shot examples, resulting in input lengths of approximately 4K, 8K, and 16K tokens respectively. The few-shot data were sourced from the GSM8K training set. Since the test set does not overlap with the training set, the answers remain undisclosed to the LLM.
Unlike long-context tasks LongBench, CoT mathematical reasoning tasks present distinct challenges. For example, the retrieval-based SOTA method, Quest, excels on long-context benchmarks but performs poorly on CoT tasks, showing a 16.91% accuracy drop at a sequence length of 16K. In contrast, AttentionPredictor achieves a significantly smaller accuracy loss of just 0.91%. Moreover, across all sequence lengths, our method outperforms baselines in most cases.
Table 7: Evaluation results on the long n-shot CoT GSM8K task with different input context lengths with Llama-3.1-8B-Instruct.
GSM8K+LLaMA-3.1
Method Budget Prompt length
4K 8K 16K Average
Full cache Full 56.79 55.27 54.13 55.40
StreamingLLM 1K 54.74 49.96 50.94 51.88
H2O+ 1K 57.16 52.01 52.16 53.78
Quest 1K 48.52 45.26 37.22 43.67
SnapKV 1K 53.45 48.67 49.20 50.44
AttentionPredictor 1K 56.48 53.30 53.22 54.33
Table 8: The detailed evaluation results from LongBench of all tasks. Since the official pre-trained model of SeerAttention for LongChat is unavailable, we only evaluate its performance on LLaMA-3.1-8B-Instruct.
@@ Appendix C Additional Experimental Results on Accuracy
4096 AttentionPredictor 28.67 47.07 55.87 57.46 48.35 27.91 33.12 24.76 26.62 73.50 90.30 43.40 6.75 100.00 58.03 48.85 48.17
C.2 Additional Results on Longbench Benchmark
SeerAttention utilizes a learnable module to aid cache compression. We use the officially released SeerAttention model for our tests. Since the official pre-trained model of SeerAttention for LongChat is unavailable, we evaluate its performance on LLaMA-3.1-8B-Instruct.
Table 8 indicates that our method outperforms SeerAttention, which also employs a learnable module. In particular, with a cache budget of 512, our method achieves an average accuracy 7.27% higher than SeerAttention. This could be attributed to SeerAttention encoding keys and queries without directly modeling attention scores, thus failing to accurately retrieve important token blocks under extreme resource constraints. Moreover, by employing a single unified model for all LLM layers, our approach results in a total model size that is merely 0.02% of SeerAttention’s, further attesting to the effectiveness of our method.
C.3 Results on InfiniteBench
To evaluate the effectiveness of our proposed method, we conduct experiments on selected tasks from the InfiniBench [46] benchmark, an extremely long-context benchmark with an average context length 214K. We use LLaMA-3-8B as the base model and focus on the Math Find and Choice tasks, which require numerical reasoning and multi-option selection, respectively. We compare our approach against several state-of-the-art KV compression baselines under an 8K token KV cache budget.
As shown in Table 9, AttentionPredictor achieves the highest average performance (44.9), outperforming Quest (43.4), SnapKV (42.5), and InfLLM (33.7). Notably, on the Math Find task, which is particularly sensitive to the preservation of precise attention information, our method attains 34.3—significantly higher than all baselines. These results suggest that AttentionPredictor is more effective at identifying and preserving the most relevant attention information under constrained memory, enabling better retention of long-range dependencies.
Table 9: Evaluation on InfiniteBench.
InfiniteBench with Llama-3-8B
KV Budget Method Math Find Choice Average
Full Full cache 21.7 44.1 32.9
8K InfLLM [32] 23.7 43.7 33.7
8K SnapKV [13] 27.4 57.6 42.5
8K Quest [31] 32.3 54.5 43.4
8K AttentionPredictor 34.3 55.5 44.9
C.4 Results on RULER QA Benchmark
We further evaluate our method on the RULER QA benchmark [47], which is designed to assess long-context understanding under varying context lengths and constrained KV cache budgets. The results are summarized in Table 10. The results demonstrate that with a highly constrained budget of only 1K tokens, our AttentionPredictor consistently and significantly outperforms the SnapKV baseline across all tested context lengths.
Table 10: Evaluation results on the RULER QA benchmark with different input context lengths with Llama-3.1-8B-Instruct.
RULER QA Benchmark with Llama-3.1-8B-Instruct
Context Length
Method KV Budget   4K   8K   16K
Full cache Full  75.0  73.0  69.0
Snap KV 1K  63.7  71.7  68.0
AttentionPredictor 1K  73.4  71.9  68.8
C.5 Results on MMLU Benchmark
To further assess the effectiveness of our approach, we conduct experiments on the MMLU benchmark [49]. Since the default evaluation mode in the official lm-eval library is a single-step prediction task, it does not fully reflect the benefits of our method, which is designed to optimize multi-step decoding. Therefore, we adopt the cot_fewshot configuration, which applies an n-shot Chain-of-Thought (CoT) prompting strategy and involves a multi-step reasoning process. This setup provides a more suitable testbed for evaluating our decoding-stage optimizations.
As shown in Table 11, our proposed AttentionPredictor achieves performance very close to that of the full-cache model while consistently outperforming the SnapKV baseline under different KV cache budgets. These results demonstrate that our method remains effective even on relatively short input tasks, provided that the decoding process requires multi-step reasoning.
Table 11: Evaluation results on the MMLU benchmark with few-shot CoT.
MMLU Benchmark with Llama-3.1-8B-Instruct
Method Budget
512 1024
Full cache 67.17 67.17
Snap KV 65.60 66.17
AttentionPredictor 66.27 66.64
C.6 Results on GPQA Benchmark
To further validate our approach on expert-level tasks, we also tested it on the GPQA benchmark [50]. GPQA is a challenging dataset of graduate-level multiple-choice questions curated by domain experts across biology, physics, chemistry, and philosophy. It is similar in format to MMLU but significantly more difficult, requiring deep expert-level reasoning.
As shown in Table 12, our method again proves effective, maintaining high accuracy in a scenario requiring deep reasoning.
Table 12: Evaluation results on the GPQA benchmark with n-shot CoT.
Method Budget 5-shot 10-shot 15-shot Average
Llama-3.1-8B-Instruct Full 31.31 37.88 31.31 33.50
SnapKV 1K 22.73 29.80 23.23 25.25
AttentionPredictor 1K 31.31 33.33 31.82 32.15
C.7 Prediction Model
Table 13: Prediction accuracy and parameter numbers of different model implementations.
LongBench+Longchat
Model type Total params Recovery rate
MLP 49K 92.88
LSTM-block 3.2M 91.75
CNN-block 21K 91.94
CNN 4.9K 93.65
Our prediction task can be framed as modeling spatio-temporal information inherent in the token sequences, a problem for which various architectures have been explored [43, 44, 57, 58]. We evaluate the performance of various prediction model implementations. MLP is applied to the sequence dimension to handle variations in token length. For LSTM, attention is divided into 16 width blocks and predictions are independent for each block, referred to as LSTM-block in Table 13. Similarly, CNN-block utilizes a block-wise prediction approach, employing a fully connected layer for fixed block lengths. Finally, the CNN model, as the primary setting in our experiments, predicts all attention scores simultaneously. As reported in Table 13, CNN outperforms other models, achieving the highest attention prediction accuracy by effectively capturing attention patterns. MLP failed to account for neighboring token interactions, while LSTM-block and CNN-block were restricted to block-level information without global context. Notably, CNN also required the fewest parameters and was the most memory-efficient during inference.
C.8 Ablation Study Results of Hyperparameters
@@
We evaluate the impact of hyperparameters on our method. Specifically, we train the predictor using the default parameters of the main experiment, i.e., $b=16$, $H=64$. We employ full attention without sparse computation during training, which is equivalent to performing distribution error calibration at each step, i.e., $M=1$. Subsequently, we test the performance under different hyperparameters using the trained model. Other parameters during testing remain consistent with the main experiment, with a KV budget of 1K. We evaluate six representative tasks in Longbench. The results are shown in Table 14.
Calibration step $M$. Table 14 shows the average performance of AttentionPredictor with calibration steps ranging from 1 to 20. Performance improves with shorter calibration intervals, indicating that the calibration scheme reduces cumulative errors caused by differences between sparse and original attention distributions. However, higher calibration frequencies increase computational costs, creating a trade-off between accuracy and efficiency.
History step $H$. As depicted in Table 14, we evaluate the average accuracy of AttentionPredictor over a range of history steps. Overall, the performance gap compared to the full cache remains consistently small. Notably, the middle value of $H$ maximizes performance by providing the necessary information for pattern recognition, while mitigating redundancy caused by the decaying self-correlation of attention scores.
@@
C.9 Necessity of the Learnable Attention Predictor
A pertinent question arises regarding the necessity of a learnable attention predictor: given the potential for high similarity between attention distributions of adjacent tokens or layers, could one simply utilize the attention map from a preceding token or layer as a direct estimate? To address this, we conducted a comparative experiment evaluating our proposed AttentionPredictor against two such heuristic baselines: (1) Previous-Layer Attention, which uses the attention map from the corresponding position in the immediately preceding layer, and (2) Previous-Token Attention, which employs the attention map of the immediately preceding token within the same layer.
Table 15: Performance comparison on LongBench with Llama3.1-8B-Instruct.
Budget Method
NrtvQA
Qasper
MF-en
HotpotQA
2WikiMQA
Musique
GovReport
QMSum
MultiNews
TREC
TriviaQA
SAMSum
PCount
PRe
Lcc
RB-P
Average↑
Full Full cache 29.28 45.36 56.12 56.67 49.90 32.43 33.77 25.20 27.08 74.00 91.48 40.85 5.07 99.50 56.86 48.15 48.17
1024 Previous layer 26.04 33.93 46.17 53.50 46.04 26.80 26.27 21.94 25.81 64.00 83.48 35.74 10.75 99.50 48.60 41.58 42.51
1024 Previous token 28.90 39.84 55.37 55.63 48.66 30.30 27.36 23.45 25.73 69.50 88.95 43.01 8.50 99.56 56.96 47.40 46.82
1024 AttentionPredictor 28.50 47.35 54.99 57.53 45.91 31.63 32.52 24.51 26.78 73.00 89.54 42.12 6.49 99.56 56.96 47.43 47.80
The results of this evaluation, performed on the LongBench benchmark using Llama3.1-8B-Instruct, are presented in Table 15. Employing previous-layer attention yields an average score of 42.51%, which is a substantial 5.29% lower than our AttentionPredictor’s average of 47.80%. This performance decrement underscores the limitation of direct cross-layer attention transference, primarily because attention patterns often exhibit considerable variation across different layers, making the previous layer’s attention a less reliable proxy for identifying critical tokens in the current layer.
The previous-token attention strategy achieves a higher average score of 46.82%. This is 4.31% better than the previous-layer approach and aligns with the observation that attention mechanisms frequently display sequential patterns along the token dimension within the same layer. However, despite this improvement, relying solely on the single immediately preceding token is still suboptimal. Our learnable AttentionPredictor, with an average score of 47.80%, outperforms the previous-token attention baseline by 0.98%. This superiority stems from our model’s capacity to learn and leverage more complex, dynamic attention patterns from several previous tokens, rather than being confined to the heuristic of the single last token, which may not adequately capture evolving contextual dependencies.
Therefore, these comparative results affirm that while attention similarities exist, simplistic heuristic estimations are insufficient. The empirical evidence highlights the necessity and distinct advantage of our proposed learnable AttentionPredictor for achieving more accurate anticipation of attention distributions.
C.10 Comparison between MInference and AttentionPredictor
To further evaluate the effectiveness of our proposed AttentionPredictor, we compare its performance against MInference [16]. MInference operates by categorizing attention heads into three predefined sparse patterns (A-shape, Vertical-Slash, and Block-Sparse) and uses a kernel-aware search to assign the optimal pattern to each head offline. During inference, it dynamically builds sparse indices based on the assigned pattern and the specific input.
We consider two variants of MInference for a comprehensive comparison: (1) MInference (prefill): The standard version where the pattern-fitting algorithm is run only during the pre-filling stage. (2) MInference (decode): A stronger, more dynamic variant where the computationally expensive pattern-fitting algorithm is re-run at every decoding step to adapt to new contexts.
As shown in Table 16, AttentionPredictor consistently outperforms both MInference variants across all tasks, achieving an average accuracy of 94.25%. This demonstrates that by learning from recent history instead of relying on fixed pattern templates, AttentionPredictor more accurately captures the dynamic nature of attention during decoding.
Table 16: Attention prediction accuracy (%) on Llama-3.1-8B-Instruct with 1K KV budget.
Method QA Summary Math Average
MInference (prefill) 90.51 84.45 93.47 89.48
MInference (decode) 92.18 88.74 94.67 91.86
AttentionPredictor 94.95 91.72 96.10 94.25
C.11 Top-K Selection Strategy Analysis
Why Top-K Outperforms Sampling? While prior works such as MagicPIG [59] suggest that Top-K may introduce bias in certain tasks, our experiments reveal a critical insight: prediction accuracy matters more than the design of the selection policy. As shown in Table 17, we compare the performance of our AttentionPredictor and MagicPIG on the LongBench benchmark using Llama-3.1-8B-Instruct (MagicPIG’s results are taken from the original paper). The results demonstrate that when combined with our high-precision AttentionPredictor, even a simple Top-K strategy achieves superior performance compared to MagicPIG’s more sophisticated sampling approach. This indicates that high-quality attention predictions can enable basic selection policies to outperform more complex alternatives. Furthermore, it is important to note that our AttentionPredictor is decoupled from the selection policy and can be flexibly integrated with various methods, including both Top-K and importance sampling.
Table 17: Longbench task results on Llama-3.1-8B-Instruct with 10% KV budgets.
Method Qasper RB-P Lcc Pre TREC TriviaQA Average
Full Cache 45.36 48.15 56.86 99.50 74.00 91.48 69.22
MagicPig 43.96 46.45 57.06 100.00 74.40 91.38 68.87
AttentionPredictor 45.50 50.26 59.41 99.67 74.00 88.61 69.58
Appendix D Additional Experimental Results on Efficiency
D.1 Training Efficiency
We evaluate the impact of the training data ratio on prediction accuracy on the hotpotqa dataset of Longbench. The predictor’s performance improved only slightly, from 96.0% to 97.5%, as the proportion of training samples increased from 3% to 70%. This efficiency can be attributed to two factors. First, the predictor only needs to capture the relative patterns rather than precisely predict attention scores, which is easier. Second, since attention patterns are inherent characteristics of the LLM, the predictor exhibits strong generalization capabilities. These findings suggest that a relatively small amount of training data is adequate for the predictor to effectively learn the underlying attention dynamics.
Table 18: Accuracy of predictor with different training sample ratios. The test set is a 20% split from the datasets.
Training Samples Ratio 0.03 0.05 0.1 0.2 0.5 0.7
Accuracy (%) on Test Samples 96.0 96.2 95.9 96.6 97.2 97.5
D.2 Computational Overhead of AttentionPredictor.
We have quantified the inference overhead of our predictor model. As shown in Table 19, both the latency and memory footprint are minimal, ensuring that the prediction process does not become a bottleneck.
Table 19: The computational overhead of the prediction model.
Context Length Prediction Latency (ms) Peak Memory Usage (MB)
4K 0.28 97
8K 0.47 194
16K 0.91 388
32K 2.22 776
D.3 Memory Efficiency of Cross-Token Prefetching
We have evaluated the memory footprint of our proposed cross-token prefetching framework during decoding. As shown in Table 20, the framework consistently consumes less GPU memory compared to the standard cross-layer approach. This advantage arises because the memory savings from loading only a small, sparse KV cache outweighs the potential overhead from the cross-token prefetching mechanism.
Table 20: Peak GPU memory consumption (GB) of LLaMA-3.1-8B during decoding.
Context Length Standard Cross-Layer Our Cross-Token Memory Reduction
4K 32.9 30.1 8%
8K 35.9 30.3 16%
16K 41.9 30.5 27%
32K 53.9 31.1 42%
Appendix E Limitations
Our current investigation is primarily centered on a specific range of LLM architectures and sizes, and the generalizability of observed attention patterns to vastly different model types warrants further study. The necessary conditions for a specific attention pattern can be explored in the future work.
Appendix F Boarder Impacts
Academic Impact: Our study of LLM attention score patterns and their predictability may offer insights into attention dynamics, potentially informing future research on these characteristics. By exploring attention score patterns, this work could also be relevant to advancements in KV cache compression, possibly contributing to the development of more memory-efficient model designs.
Societal Impact: Through the potential implications for KV cache compression via attention predictability, this research could contribute to more efficient LLM inference. This might lead to reduced latency in AI applications and lower energy consumption. Such improvements could potentially enhance the accessibility of LLMs and support more sustainable AI practices.
Appendix G License
We include the following licenses for the code, models, and benchmarks we used in this paper.
Benchmarks: LongBench [45]: License, AIME 2024 [48]: License, GSM8K [60]: License, Needle In A Haystack [51]: License, InfiniteBench [46]: License. Models: LLaMA-3.1-8B-Instruct [54]: License, LongChat [53]: License. Code: H2O [11]: License, Minference [16]:License, SeerAttention [14]: License.